# A characterization of the local structure of two-dimensional sets with positive reach

### v1

@@ 5.3. Some properties of skewered sets with positive reach
To prove (ii), first observe that since $A$ is compact, $\rho(A)$ has the same property by (i). To prove ${\rm reach}\,\rho(A)>0$, we first observe that by Lemma 5.12 (vi), for each $\lambda\in\Lambda^{\prime}$, ${\rm reach}\,\overline{A}_{\lambda}>r$. Let $R_{\lambda}$ be the $2$-rotation agreeing with $\rho$ on $\pi^{-1}(\lambda)$, $\lambda\in\Lambda^{\prime}$. Then for each $\lambda\in\Lambda^{\prime}$, we have $\rho(\overline{A}_{\lambda})=R_{\lambda}(\overline{A}_{\lambda})$ by Lemma 5.12 (v) and, hence, ${\rm reach}\,\rho(\overline{A}_{\lambda})={\rm reach}\,\overline{A}_{\lambda}>r$. (The last equality follows from the easy fact that isometries in $\mathbb{R}^{d}$ preserve the reach.) Now we will apply Lemma 2.19. So consider two arbitrary points $x,y\in\rho(A)$; we can suppose that $\pi(x)\leq\pi(y)$. We will distinguish two cases.
a) There exists $\lambda\in\Lambda^{\prime}$ such that $\pi(x),\pi(y)\in\overline{\lambda}$. Since $x,y\in\rho(\overline{A}_{\lambda})$, Lemma 2.19 and Remark 2.20 give that
(51) $\left|\frac{x+y}{2}-s\right|\leq\frac{|x-y|^{2}}{4r}\ \text{ for some }\ s\in\rho(\overline{A}_{\lambda})\subset\rho(A).$
@@
(52) $\hskip-25.60747pt\leq\tfrac{F}{2}|\pi(x)-z|^{2}+\tfrac{F}{2}|\pi(y)-z|^{2}\leq F|\pi(x)-\pi(y)|^{2}\leq F|x-y|^{2}.$
Using (51), (52), Lemma 2.19 and Remark 2.20, we easily obtain that ${\rm reach}\,\rho(A)>0$.
@@
This easily implies that $\operatorname{Tan}(\rho(A),z)\subset\mathbb{R}$, and, since $p,q]\subset\rho(A)$, we get $z\in T_{1}(\rho(A))$. Thus $T_{1}(A)=T_{1}(\rho(A))$ follows. ∎
In dimension 2, skewered sets with positive reach are closely connected to $\operatorname{B}$-sets.
@@
Let $M\subset\mathbb{R}^{2}$ be a set with positive reach skewered on a segment $c,d]\subset\mathbb{R}$, $c<0<d$, and let $0\in T_{1}(M)$. Then there exists a $\operatorname{B}_{+}$-set $B\subset\mathbb{R}^{2}$ such that $(M,0)\sim(B,0)$.
@@
Since $M\subset\mathbb{R}^{2}$, ${\rm reach}\,(M,0)>0$ and $0\in T_{1}^{+}(M)$, we can apply [23, Theorem 6.4] and get that $M$ is of type $T_{2}$ at 0 which means, by definition, that there exists a linear isometry $G:\mathbb{R}^{2}\to\mathbb{R}^{2}$ such that $G(M\cap B(0,r))$ is a $\tilde{T}_{r}^{2}$-set for some $r>0$ (see [23, Definition 6.3]). Comparing the tangent cone $\operatorname{Tan}(M,0)=W$ with that of a $\tilde{T}_{r}^{2}$-set at 0, we see that the linear isometry $G$ maps $e_{1}$ onto $\pm e_{1}$. Thus, the matrix of $G$ is diagonal with $\pm 1$ on the main diagonal, i.e., $G=I$ (identity), or $G:(x,y)\mapsto(-x,y)$, or $G:(x,y)\mapsto(x,-y)$, or $G:(x,y)\mapsto(-x,-y)$.
As $M$ is skewered on $[c,d]$, it contains a neighbourhood of the origin in $W$, and the $\tilde{T}_{r}^{2}$-set $G(M\cap B(0,r))$ has the same property. It is easy to see that a $\tilde{T}_{r}^{2}$-set containing a neighbourhood of the origin in $W$ locally agrees at 0 with a $B_{+}$-set. Consequently, there exists a $\operatorname{B}_{+}$-set $B^{\prime}\subset\mathbb{R}^{2}$ such that $(G(M),0)\sim(B^{\prime},0)$. Due to the special form of the isometry $G$, and using Remark 2.7, we obtain that the preimage $B:=G^{-1}(B^{\prime})$ is again a $B_{+}$-set. Since clearly $(M,0)\sim(B,0)$, the proof is complete. ∎
@@
Choose $0<2\omega<r<{\rm reach}\,S$ and denote $S^{*}:=S\cap\overline{B}(0,\omega)$. Due to Lemma 5.12 (ii), ${\rm reach}\,S^{*}>r$, $\dim S^{*}\leq 2$, $S^{*}$ skewered on $c^{*},d^{*}]:=[c,d]\cap[-\omega,\omega]$, and $0\in T_{1}(S^{*})$. Using also Lemma 5.11 we see that each leave of $S^{*}$ is a subset of a leave of $S$, hence $S^{*}$ has planar leaves as well. Let $\tilde{\rho}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the multirotation associated with $T_{1}(S^{*})$ (see Definition 1.2) defined as follows (recall Notation 5.5): If $\lambda$ is a component of $W\setminus T_{1}(S^{*})$ with $\lambda\cap[c^{*},d^{*}]\neq\emptyset$ and $V_{\lambda}\in G(d,2)$ is the plane containing the leave $S^{*}_{\lambda}$, choose a unit vector $v_{\lambda}\in V_{\lambda}\cap W^{\perp}$ and define the $2$-rotation $\tilde{R}_{\lambda}$ as identity if $v_{\lambda}=\pm e_{2}$ and as the $2$-rotation mapping $v_{\lambda}$ to $e_{2}$ and being identity on $\operatorname{span}\{v_{\lambda},e_{2}\}^{\perp}$ otherwise. If $\lambda\cap[c^{*},d^{*}]=\emptyset$ we set $\tilde{R}_{\lambda}:={\rm id}$. The multirotation $\tilde{\rho}$ is then determined by the family $(\tilde{R}_{\lambda})$, where $\lambda$ are the components of $W\setminus T_{1}(S^{*})$.
@@
Note that $B^{\prime}$ has positive reach also as a subset of $\mathbb{R}^{2}$ (see Remark 2.14). By Lemma 5.16, there exists a $\operatorname{B}_{+}$-set $B\subset\mathbb{R}^{2}$ and $0<\tau<\min\{{\rm reach}\,B,{\rm reach}\,B^{\prime}\}$ such that
@@
Note that $\tilde{\rho}^{-1}$ is a multirotation associated with $T_{1}(B^{\prime})=T_{1}(S^{*})$. Since $T_{1}(B^{\prime})\cap\overline{B}(0,\tau)=T_{1}(B)\cap\overline{B}(0,\tau)$ by (53) and Lemma 5.11 (applied to both $B$ and $B^{\prime}$), there exists a multirotation $\rho$ associated with $T_{1}(B)$ such that $\rho|_{\pi^{-1}([-\tau,\tau])}=\tilde{\rho}^{-1}|_{\pi^{-1}([-\tau,\tau])}$, see Lemma 5.4 and, consequently, $S^{*}\cap B(0,\tau)=\rho(B)\cap B(0,\tau)$. Hence $(S,0)\approx(S^{*},0)\approx(\rho(B),0)$ and the assertion follows. ∎
@@
If $a\in T_{1}^{+}(A)$, the first statement (which is equivalent to Proposition 5.3, see Subsection 5.1) follows from Propositions 5.8, 5.9 and 5.10.
Let us consider now the case $a\in T_{1}^{-}(A)$. Choose $0<r<{\rm reach}\,A$ and let $u$ be the unit vector such that $\operatorname{Tan}(A,a)=\{tu:\,t\geq 0\}$. Denoting $S:=\{a+tu:\,-\frac{r}{4}\leq t<0\}$ and $A^{*}:=A\cup S$, we have clearly $a\in T_{1}^{+}(A^{*})$ and [23, Lemma 3.8] gives ${\rm reach}\,A^{*}\geq r/4$. We apply the already proved assertion and find a $\operatorname{B}_{+}$-set $B^{*}$ and a multirotation $\rho^{*}$ associated with $T_{1}(B^{*})$ such that $(A^{*},a)\approx(\rho^{*}(B^{*}),0)$. By Remark 5.2 (i) there exist a $C^{1,1}$-diffeomorphism $\Phi$ defined on a neighbourhood $U$ of $a$ and $\varepsilon>0$ such that
@@
$\Phi(S_{\delta})=\rho^{*}(B^{*})\cap\{(x,y):\,x<0\}\cap B(0,\varepsilon^{\prime}).$
@@
$=\rho^{*}(B^{*})\cap\{(x,y):\,x\geq 0\}\cap B(0,\varepsilon^{\prime})$
$=\rho^{*}(B^{*}\cap\{(x,y):\,x\geq 0\})\cap B(0,\varepsilon^{\prime}).$
Note that $B:=B^{*}\cap\{(x,y):\,x\geq 0\}$ is a $\operatorname{B}_{-}$-set. Using Lemma 5.4, we can find a multirotation $\rho$ associated with $T_{1}(B)$ and such that $\rho(B)=\rho^{*}(B)$. Then
@@
which proves Proposition 5.3 (and, thus, also the first statement of Theorem 1.3) in the case $a\in T_{1}^{-}(A)$.
For the second statement of Theorem 1.3, assume that $B\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ is a $\operatorname{B}$-set, $\rho$ a multirotation associated with $T_{1}(B)$, $U$ an open neighbourhood of $\rho(B)$ and $\Phi:U\to\mathbb{R}^{d}$ a $C^{1,1}$-diffeomorphism. Then $B$ has positive reach by Lemma 2.22 and it is clearly compact, hence $\rho(B)$ is compact and has positive reach by Lemma 5.15 (ii). Setting $A:=\Phi(\rho(B))$, we have ${\rm reach}\,A>0$ by Lemma 2.21 and $a\in T_{1}(A)$ by Lemma 5.15 (iii) and Lemma 3.2 (iii). This implies the second assertion of Theorem 1.3 and completes the proof. ∎
@@
Let $A\subset\mathbb{R}^{d}$ be a two-dimensional set with positive reach and $a\in A$. If $a\in T_{0}(A)$ then $a$ is an isolated point in $A$ and the assertion is obvious. If $a\in T_{1}(A)$ then we apply Theorem 1.3 and obtain a $\operatorname{B}$-set $B\subset\mathbb{R}^{2}$, a multirotation $\rho$ associated with $T_{1}(B)$, a neighbourhood $U$ of $\rho(B)$ in $\mathbb{R}^{d}$ and a $C^{1,1}$-diffeomorphism $\Phi:U\to\mathbb{R}^{d}$ such that $\Phi(\rho(B))$ is a neighbourhood of $a$ in $A$. The set $\rho(B)$ is compact by Lemma 5.15 (ii), hence $\Phi(\rho(B))$ is compact as well. Since both $\Phi$ and $\Phi^{-1}$ are locally Lipschitz (see Remark 2.1 (i)), their restrictions to compact sets $\rho(B)$ and $\Phi(\rho(B))$ (respectively) are Lipschitz by [16, Proposition A48], i.e., $\Phi$ is bi-Lipschitz on $\rho(B)$. Since $\rho$ is bi-Lipschitz on $B$ by Lemma 5.15 (i), we can choose $\Psi:=\Phi\circ\rho|_{B}$ and the assertion follows.

### v2

@@ 5.3. Some properties of skewered sets with positive reach
To prove (ii), first observe that since $A$ is compact, $\rho(A)$ has the same property by (i). To prove ${\rm reach}\,\rho(A)>0$, we first observe that by Lemma 5.12 (vi), for each $\lambda\in\Lambda^{\prime}$, ${\rm reach}\,\overline{A}_{\lambda}>r$. Let $R_{\lambda}$ be the $2$-rotation agreeing with $\rho$ on $\pi^{-1}(\lambda)$, $\lambda\in\Lambda^{\prime}$. Then for each $\lambda\in\Lambda^{\prime}$, we have $\rho(\overline{A}_{\lambda})=R_{\lambda}(\overline{A}_{\lambda})$ by Lemma 5.12 (v) and, hence, ${\rm reach}\,\rho(\overline{A}_{\lambda})={\rm reach}\,\overline{A}_{\lambda}>r$. (The last equality follows from the easy fact that isometries in $\mathbb{R}^{d}$ preserve the reach.) Now we will infer from Lemma 2.20 that ${\rm reach}\,\rho(A)\geq r^{\prime}:=\min\{\tfrac{r}{2},\tfrac{1}{8F}\}.$ To this end, consider two arbitrary points $x,y\in\rho(A)$ with $|x-y|<2r^{\prime}$; we can suppose that $\pi(x)\leq\pi(y)$. We will distinguish two cases.
a) There exists $\lambda\in\Lambda^{\prime}$ such that $\pi(x),\pi(y)\in\overline{\lambda}$. Since $x,y\in\rho(\overline{A}_{\lambda})$, ${\rm reach}\,\overline{A}_{\lambda}>r$ and $r^{\prime}\leq r/2$, Lemma 2.20 and Remark 2.21 give that
(51) $\left|\frac{x+y}{2}-s\right|\leq\frac{|x-y|^{2}}{4r}\leq\frac{|x-y|^{2}}{8r^{\prime}}\ \text{ for some }\ s\in\rho(\overline{A}_{\lambda})\subset\rho(A).$
@@
$\leq\tfrac{F}{2}|\pi(x)-z|^{2}+\tfrac{F}{2}|\pi(y)-z|^{2}\leq F|\pi(x)-\pi(y)|^{2}$
(52) $\leq F|x-y|^{2}\leq\frac{|x-y|^{2}}{8r^{\prime}}.$
Using (51), (52), Lemma 2.20 and Remark 2.21, we easily obtain that ${\rm reach}\,\rho(A)\geq r^{\prime}>0$.
@@
This easily implies that $\operatorname{Tan}(\rho(A),z)\subset\mathbb{R}$, and, since $z\in[p,q]\subset\rho(A)$, we get $z^{*}=z\in T_{1}(\rho(A))$. Thus $T_{1}(A)=T_{1}(\rho(A))$ follows. ∎
In $\mathbb{R}^{2}$, skewered sets with positive reach are closely connected to $\operatorname{B}$-sets.
@@
Let $M\subset\mathbb{R}^{2}$ be a set with positive reach skewered on a segment $c,d]\subset\mathbb{R}$, $c<0<d$, and let $0\in T_{1}(M)$. Then there exists a $\operatorname{B}_{+}$-set $B\subset\mathbb{R}^{2}$ such that $M\sim_{0}B$.
@@
Since $M\subset\mathbb{R}^{2}$, ${\rm reach}\,(M,0)>0$ and $0\in T_{1}^{+}(M)$, we can apply [22, Theorem 6.4] and get that $M$ is of type $T_{2}$ at 0 which means, by definition, that there exists a linear isometry $G:\mathbb{R}^{2}\to\mathbb{R}^{2}$ such that $G(M\cap B(0,r))$ is a $\tilde{T}_{r}^{2}$-set for some $r>0$ (see [22, Definition 6.3]). Comparing the tangent cone $\operatorname{Tan}(M,0)=W$ with that of a $\tilde{T}_{r}^{2}$-set at 0, we see that the linear isometry $G$ maps $e_{1}$ onto $\pm e_{1}$. Thus, the matrix of $G$ is diagonal with $\pm 1$ on the main diagonal, i.e., $G=I$ (identity), or $G:(x,y)\mapsto(-x,y)$, or $G:(x,y)\mapsto(x,-y)$, or $G:(x,y)\mapsto(-x,-y)$.
As $M$ is skewered on $[c,d]$, it contains a neighbourhood of 0 in $W$, and hence the $\tilde{T}_{r}^{2}$-set $G(M\cap B(0,r))$ has the same property. It is easy to see that a $\tilde{T}_{r}^{2}$-set containing a neighbourhood of the origin in $W$ locally agrees at 0 with a $B_{+}$-set. Consequently, there exists a $\operatorname{B}_{+}$-set $B^{\prime}\subset\mathbb{R}^{2}$ such that $G(M)\sim_{0}B^{\prime}$. Due to the special form of the isometry $G$, and using Remark 2.8, we obtain that the preimage $B:=G^{-1}(B^{\prime})$ is again a $B_{+}$-set. Since clearly $M\sim_{0}B$, the proof is complete. ∎
@@
Choose $\omega,r$ with $0<2\omega<r<{\rm reach}\,S$ and denote $S^{*}:=S\cap\overline{B}(0,\omega)$. Due to Lemma 5.12 (ii), ${\rm reach}\,S^{*}>r$, $\dim S^{*}\leq 2$, $S^{*}$ skewered on $c^{*},d^{*}]:=[c,d]\cap[-\omega,\omega]$, and $0\in T_{1}(S^{*})$. Using also Lemma 5.11 we see that each leave of $S^{*}$ is a subset of a leave of $S$, hence $S^{*}$ has planar leaves as well. Let $\tilde{\rho}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the multirotation associated with $T_{1}(S^{*})$ (see Definition 1.2) defined as follows (recall Notation 5.5): If $\lambda$ is a component of $W\setminus T_{1}(S^{*})$ with $\lambda\cap[c^{*},d^{*}]\neq\emptyset$ and $V_{\lambda}\in G(d,2)$ is the plane containing the leave $S^{*}_{\lambda}$, choose a unit vector $v_{\lambda}\in V_{\lambda}\cap W^{\perp}$ and define the $2$-rotation $\tilde{R}_{\lambda}$ as identity if $v_{\lambda}=\pm e_{2}$ and as the $2$-rotation mapping $v_{\lambda}$ to $e_{2}$ and being identity on $\operatorname{span}\{v_{\lambda},e_{2}\}^{\perp}$ otherwise. If $\lambda\cap[c^{*},d^{*}]=\emptyset$ we set $\tilde{R}_{\lambda}:={\rm id}$. The multirotation $\tilde{\rho}$ is then determined by the family $(\tilde{R}_{\lambda})$, where $\lambda$ are the components of $W\setminus T_{1}(S^{*})$.
@@
Note that $B^{\prime}$ has positive reach also as a subset of $\mathbb{R}^{2}$ (see Remark 2.15). By Lemma 5.16, there exists a $\operatorname{B}_{+}$-set $B\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ and a number $0<\tau<\min\{{\rm reach}\,B,{\rm reach}\,B^{\prime}\}$ such that
@@
Note that $\tilde{\rho}^{-1}$ is a multirotation associated with $T_{1}(B^{\prime})=T_{1}(S^{*})$. Since $T_{1}(B^{\prime})\cap\overline{B}(0,\tau)=T_{1}(B)\cap\overline{B}(0,\tau)$ by (53) and Lemma 5.11 (applied to both $B$ and $B^{\prime}$), there exists a multirotation $\rho$ associated with $T_{1}(B)$ such that $\rho|_{\pi^{-1}([-\tau,\tau])}=\tilde{\rho}^{-1}|_{\pi^{-1}([-\tau,\tau])}$ (see Lemma 5.4) and consequently $S^{*}\cap B(0,\tau)=\rho(B)\cap B(0,\tau)$. Hence $S\sim_{0}S^{*}\sim_{0}\rho(B)$ and the assertion follows. ∎
@@
If $a\in T_{1}^{+}(A)$, the first part of the theorem (which is equivalent to Proposition 5.3, see Subsection 5.1) follows from Propositions 5.8, 5.9 and 5.10. So it is sufficient to prove Proposition 5.3 in the case $a\in T_{1}^{-}(A)$.
In this case, choose $0<r<{\rm reach}\,A$ and let $u$ be the unit vector such that $\operatorname{Tan}(A,a)=\{tu:\,t\geq 0\}$. Denoting $S:=\{a+tu:\,-\frac{r}{4}\leq t<0\}$ and $A^{*}:=A\cup S$, we have ${\rm reach}\,A^{*}\geq r/4$ by [22, Lemma 3.8], and clearly $a\in T_{1}^{+}(A^{*})$. We apply the already proved assertion and find a $\operatorname{B}_{+}$-set $B^{*}\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ and a multirotation $\rho^{*}$ associated with $T_{1}(B^{*})$ such that $(A^{*},a)\approx(\rho^{*}(B^{*}),0)$. By Remark 5.2 (i) there exist a $C^{1,1}$-diffeomorphism $\Phi$ defined on a neighbourhood $U$ of $a$ and $\varepsilon>0$ such that
@@
$\Phi(S_{\delta})=(-\varepsilon^{\prime},0)=\rho^{*}(B^{*})\cap\{(x,z):\,x<0\}\cap B(0,\varepsilon^{\prime}).$
@@
$=\rho^{*}(B^{*})\cap\{(x,z):\,x\geq 0\}\cap B(0,\varepsilon^{\prime})$
$=\rho^{*}(B^{*}\cap\{(x,z):\,x\geq 0\})\cap B(0,\varepsilon^{\prime}).$
Note that $B:=B^{*}\cap\{(x,y):\,x\geq 0\}\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ is a $\operatorname{B}_{-}$-set. Using Lemma 5.4 (applied with $\rho:=\rho^{*}$, $C:=T_{1}(B^{*})$, $C^{\prime}:=T_{1}(B)$ and $[p,q]:=[0,v]$), we can find a multirotation $\rho$ associated with $T_{1}(B)$ and such that $\rho(B)=\rho^{*}(B)$. Then
@@
which proves the assertion of Proposition 5.3.
To prove the second part of Theorem 1.3, assume that $B\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ is a $\operatorname{B}$-set, $\rho$ a multirotation associated with $T_{1}(B)$, $U\supset\rho(B)$ is open and $\Phi:U\to\mathbb{R}^{d}$ a $C^{1,1}$-diffeomorphism with $\Phi(0)=a$. Then $\rho(B)$ is compact and has positive reach by Lemma 5.15 (ii). Setting $A:=\Phi(\rho(B))$, we have ${\rm reach}\,A>0$ by Lemma 2.22. We have $0\in T_{1}(\rho(B))$ by Lemma 5.15 (iii) and, hence, $a\in T_{1}(A)$ by Lemma 3.2 (iii). Further, $\dim\rho(B)\leq 2$ since $\rho|_{B}$ is a homeomorpism by Lemma 5.15 (i) and, hence, also $\dim A\leq 2$, since $A$ is homeomorpic to $\rho(B)$. So the second part of Theorem 1.3 is proved. ∎
@@
Let $A\subset\mathbb{R}^{d}$ be a two-dimensional set with positive reach and $a\in A$. If $a\in T_{0}(A)$ then $a$ is an isolated point in $A$ and the assertion is obvious. If $a\in T_{1}(A)$ then we apply Theorem 1.3 and obtain a $\operatorname{B}$-set $B\subset\mathbb{R}^{2}$, a multirotation $\rho$ associated with $T_{1}(B)$, an open set $U\supset\rho(B)$ and a $C^{1,1}$-diffeomorphism $\Phi:U\to\mathbb{R}^{d}$ such that $\Phi(\rho(B))$ is a neighbourhood of $a$ in $A$. The set $\rho(B)$ is compact by Lemma 5.15 (ii), hence $\Phi(\rho(B))$ is compact as well. Since both $\Phi$ and $\Phi^{-1}$ are locally Lipschitz (see Remark 2.1 (i)), their restrictions to compact sets $\rho(B)$ and $\Phi(\rho(B))$ (respectively) are Lipschitz by [16, Proposition A48], i.e., $\Phi$ is bi-Lipschitz on $\rho(B)$. Since $\rho$ is bi-Lipschitz on $B$ by Lemma 5.15 (i), we can choose $\Psi:=\Phi\circ\rho|_{B}$ and the assertion follows.